PLEASE HELP!

When DNA is decoded into protein the 'letters' of the DNA sequence are 'read' in ordered triplets. For example, the sequence 'ACGTACTTA' would be read as the three ordered triplets 'ACG,' 'TAC,' and 'TTA.' Each letter can be either G, T, C, or A and each distinct triplet can code for a different protein subunit called an amino acid. How many different amino acids can be coded for?
Answer #1

The codon is three nucleotides long.

We have a total of 4 bases, A, T, G, C.

At every position of codon, we have 4 choices for a base.

So, the total number of combinations possible = 4^3 = 4 × 4 × 4 = 64.

So, by just 4 bases, we can make 64 amino acids if the codon is 3 bases long.

Please rate high.